# Range Membership Inference Attacks

### v2

@@ 5.3 Implementation details
We train 8 models on Purchase-100 and CIFAR-10, and train 4 models on CelebA and AG News. Each model is trained on half of the dataset, as described in [5, 48]. For all PyTorch models, we use Adam as our optimizer with a learning rate of 0.001. For WideResnets, we use the training code from [5]. On AG News, the models are trained for 4 epochs. On other datasets, they are trained for 100 epochs. All training is done on two Nvidia RTX 3090 GPUs. Training on AG News takes about 1 hour per epoch. Training other models takes less than one hour each.
5.4 Metrics
To evaluate the performance of RaMIA and MIA based on the same MIA algorithm, we use AUCs. The inputs to RaMIA are range queries in Table 1, while the inputs to MIA are point queries, which are range centers, in Table 1. The ground truth membership for both RaMIA and MIA is the range membership information defined in Table 2, which allows us to compare the power of both inference attack frameworks in identifying queries that are privacy leaking. In other words, both point and range TPRs/FPRs are calculated against the range membership information. Following [5], we also report (Range) TPR at small (Range) FPR for both MIA and RaMIA (Table 4).
@@ 5.4 Metrics
5.5 RaMIAs quantify privacy risks more comprehensively than MIAs
@@
As we have explained before, data points that are close enough to the training data are out of the scope of membership inference attacks. We observe from Figure 3 that range membership inference attacks are better at identifying those nearby points, and thus providing more comprehensive privacy auditing on all the four datasets we tested. We want to emphasize that the gain is remarkable if we consider how little samples were taken compared to the range sizes. On Purchase-100, there are a total of 1024 candidates, and we take less than 20% of them. On AG News, there are millions of sentences within a distance of 8. 50 sentences are too little to meaningfully cover anything in the space. Yet, limited samples can lead to noticeable gains, which further shows the current privacy quantification approach is suboptimal and needs a better framework. Due to randomness in sampling, we report the average gain of RaMIA over MIA with standard deviation in Table 3.TPRs at small FPRs are in Table 4. The improvement in AUC is summarized in Table 3.
@@
Table 3: Improvement in AUCs after switching from MIA to RaMIA across multiple iterations of random sampling. We do not randomly sample but use all available attack images in CelebA.
@@
Note that in the CelebA experiment, we use identify information as the semantic range function, making it similar to user-level inference. This further shows that RaMIA is a better and more comprehensive privacy auditing framework. In terms of algorithms, our attack strategy should dominate the simple averaging approach used in prior work [31, 23], since the trimming ratio is optimized (compared to no trimming in simple averaging). There are other user-level inference algorithms [11] that train additional shadow models and classifiers, which incurs additional computational costs. They also consider scenarios where user information is the label of their target models, e.g. facial and speech recognition systems that predict user ID as their outputs [11, 10]. This encourages their target models to explicitly cluster data based on user information, making inference easier. On the other hand, our target model, a facial attribute classifier, does not use user information anywhere in the training, making the inference harder.
Table 4: True Positive Rate under different attacks on different datasets at small false positive rates of $1\%$ and $0.1\%$. MIAs cannot be conducted on incomplete data, so we fill the missing columns with the modes and run the attack on them. Standard deviation over random sampling iterations is reported, except for CelebA, where we use all available candidates. The TPR and FPR are calculated based on the range membership information, as described in Sec 5.4. For a fair comparison, we should compare RaMIA and MIA based on the same membership testing backbone, e.g. MIA with RMIA versus RaMIA with RMIA.
@@
LOSS $0\pm 0$ $0\pm 0$ $1.17\pm 0.06$ $0.05\pm 0.03$ 1.40 0.28 $1.10\pm 0.11$ $0\pm 0$
RMIA $2.57\pm 1.58$ $0.57\pm 0.47$ $3.59\pm 0.11$ $0.80\pm 0.02$ 1.44 0.22 $0.54\pm 0.12$ $0\pm 0$
5.6 Factors affecting RaMIA performance
@@
Due to the nature of the sampling-based approach, the chance of our attack set containing a true training point scales linearly with the density of training points in the range. If we keep the sample size constant, increasing the range without including more training points in the range hurts the attack performance because the chance of the attack set including any training point gets diluted. On the other hand, if we increase the training point density, which is equivalent to increasing the probability the attacker samples a true training point, the attack performance gets boosted. Figure 4(b) shows that the performance of RaMIA increases when the range becomes larger in the CIFAR-10 experiment. Recall that the range function in CIFAR-10 is based on image augmentation methods. Increasing the range means the attacker applies more distinct augmentation methods to obtain transformed images. This increases the chance of the attacker obtaining one of the transformed versions of training images seen by the model during training, thus leading to better attack performance. In Figure 3(b), we conducted the attack assuming the attacker cannot sample any true training images. As a sanity check, we relax this assumption, and Figure 4(a) shows that RaMIA performs monotonically better when the density of training images increases from 0% to 50%, when the number of samples is constant.
(a) RaMIA on CelebA gets better when the training points available for sampling increases. The percentages are the density of training points in the distribution that the attack samples from. (Sanity Check)
(b) RaMIA performs better on CIFAR-10 when the range size increases. The range size here is equal to the number of distinct transformations applied to images.
@@
Ranges containing training points that are susceptible to MIAs are also more susceptible to RaMIAs. Researchers have previously discovered that machine learning models memorize duplicate data more [26, 6]. In our CelebA dataset, each celebrity has a different number of photos in the training set, which can be thought of as each identity having different levels of duplication in the training set. Similar to the insights from MIAs, we also observe that identities that have more training images, i.e. higher duplication rate, are more susceptible to RaMIA. Figure 6 shows the relationship between the percentile of each range’s RaMIA score within non-members’ RaMIA scores and the duplication rate. Generally speaking, identities that have more training photos are more prone to RaMIAs. Similarly, correlation can be observed on the other three datasets in our experiments, where the training points’ RaMIA score percentiles among non-members are positively correlated with their MIA score percentiles 5.
@@
5.7 Mismatched training and attack data hurts attack performance
Figure 2(c) shows that MIA underestimates the privacy risk when the augmentation used in training and attacking differs. This rings a bell as many people audit the privacy risk of image classifiers with original images, when the classifiers are often trained with a composition of augmentations. Many transformations, such as color jittering and affine transformations, always produce different final images. Other commonly used augmentation methods, such as random cropping, introduce more randomness to the pipeline. Hence, it is almost certain that the original images are never seen by the model. Therefore, we should use RaMIA for a better auditing result (Figure 3(c)).
Difference to existing augmentation-based MIAs
Existing attacks [5, 48] also use augmented queries in the attack, but with a different rationale and assumption of the attacker’s knowledge. Since they adopt the existing privacy notion based on point queries, only (augmented) images seen by the model in the training stage are considered as members. Hence, the attacker needs to know the exact train-time augmentations and augment images accordingly to not violate the privacy notion. In RaMIA, the set of augmentations is given by the challenger (Def 2), which can contain augmentations not used in training, but considered as privacy leaking. Using the aggregation method in [5] will hurt the attack performance if non-training augmentations are used. However, RaMIA is designed to be robust in this scenario (Fig 3(c)) due to its trimming process.
5.8 RaMIA on redacted data
Many large language models (LLMs) are trained with sensitive textual data. Some of the data with sensitive information redacted might be publicly available. Similar to our experiment with data with missing values, we can apply RaMIA to redacted data to identify which of them are used to train a target LLM. Accurately identifying the redacted sentences paves the way for reconstructing them as a follow-up attack. Figure 7 shows the results. In this experiment, we use spaCy [18] to mask peoples’ names to simulate the masking of personally identifiable information (PII). We then generate 10 possible sentences for each masked sentence using BERT and conduct RaMIA. The MIA performance is the average attack performance over all 10 possible sentences. The performance gap is smaller compared to that in Figure 3(d). The reason might be that BERT fails to produce diverse PII completions, making all candidate sentences similar to each other, and reducing the power of RaMIA.
@@ 5.8 RaMIA on redacted data
6 Conclusion
In this paper, we argue that membership inference attacks are only useful as a privacy audit tool when querying exact copies of training and test data. Moving the query to similar points causes a drastic decrease in performance, rendering MIAs less useful. We conclude MIAs fail to comprehensively capture the notion of privacy, and thus propose a new class of inference attack, RaMIA, that extends the notion of MIAs and covers the failure cases of MIAs by checking if a given range contains a training point. We introduce RaMIA as an attack framework that can be implemented with any existing MIA algorithm. We show that it can provide better privacy auditing with very few samples taken randomly. We hope our work can make more privacy researchers and practitioners aware of the shortcomings of MIAs, and shift their attention to RaMIAs. As it is the first paper that brings up this new framework, there is room for improvement in specific attack algorithms. For example, a better sampling process will surely increase the gap between RaMIA and MIA. Nevertheless, we have shown our framework is sensible and useful. In future work, we hope to design more powerful RaMIA strategies that are robust to the change of membership signals and datasets, especially on LLMs where we believe our privacy notion is extremely relevant.
Appendix A Attack algorithms
In this section, we explain the details of the membership inference attack algorithms used in our experiments.
LOSS
LOSS [46] computes loss values as a proxy of membership score on given points: $\text{MIA}(x;\theta)=\mathit{l}(x;\theta)$. To compute the likelihood, an easy way is to take the exponential of the negative of the loss $\text{I\kern-1.49994ptP}=\exp^{-\mathit{l}}$.
RMIA
RMIA [48] computes membership score by applying chain rule: $\text{I\kern-1.49994ptP}{(\theta|x)}=\frac{\text{I\kern-1.04996ptP}{(x|\theta)}\text{I\kern-1.04996ptP}{(\theta)}}{\text{I\kern-1.04996ptP}{(x)}}$. The score is then compared with all available population data points to obtain the percentage of population points being dominated by the given point: $\text{I\kern-1.49994ptP}_{z\in Z}{[\frac{\text{I\kern-1.04996ptP}{(\theta|x)}}{\text{I\kern-1.04996ptP}{(\theta|z)}}\geq\gamma]}$, where the term $\text{I\kern-1.49994ptP}{(\theta)}$ will cancel out with each other. The normalizing constant $\text{I\kern-1.49994ptP}{P(x)}$ is computed with reference models: $\text{I\kern-1.49994ptP}{(x)}=0.5\mathbb{E}_{\theta_{\text{IN}}}\text{I\kern-1.49994ptP}{(x|\theta_{\text{IN}})}+0.5\mathbb{E}_{\theta_{\text{OUT}}}\text{I\kern-1.49994ptP}{(x|\theta_{\text{OUT}})}$. In its offline version, the in models are unavailable. In this case, the former probabilities are approximated by the latter term $\text{I\kern-1.49994ptP}_{\text{IN}}=a\text{I\kern-1.49994ptP}_{\text{OUT}}+(1-a)$. The hyperparameter $\alpha$ is chosen based on the reference models. Specifically, one reference model is chosen as the temporary target model, and the rest are used to attack it. The value of $\alpha$ is chosen to be the best-performing value under this setting, obtained via a simple sweeping. In our experiment, we use the offline attack only. The $\alpha$ values for Purchase-100 and CIFAR-10 are taken from [48]. For CelebA, we set it to be 0.33. For AG News, we set it to be 1.0.

### v3

@@ V-C Implementation details
We train 16 models on Purchase-100, and 4 models on CelebA, CIFAR-10, and AG News. Each model is trained on half of its respective dataset, following setups from [17, 18]. For all PyTorch models, we use Adam with a learning rate of 0.001. For WideResNets, we use the training code from [17]. AG News models are trained for 4 epochs, and other datasets for 100 epochs. Training is conducted on two Nvidia RTX 3090 GPUs, with AG News taking about 1 hour per epoch and other models less than one hour each.
V-D Metrics
To evaluate the performance of RaMIA and MIA using the same membership inference backbone, we measure AUCs. The inputs to RaMIA are range queries (as specified in Table I), while the inputs to MIA are the point queries corresponding to range centers. Importantly, the ground truth membership for both RaMIA and MIA is defined according to range membership (see Table II). This means that even though the MIA attack is based on point queries, the true membership is based on whether the range contains at least one training point. If evaluated against the correct MIA privacy notion (which only considers exact training points as members), the AUC would be close to random guessing ($\approx 0.5$). Our experimental setup uses the MIAScore of the range center as a proxy to solve the range membership inference game, allowing us to compare the power of both frameworks in identifying privacy leaking queries. It also illustrates the utility of sampling within the range to construct a more meaningful score under RaMIA’s privacy notion. Following [17], we also report (Range) TPR at small (Range) FPR for both methods (see Table IV).
@@ V-D Metrics
V-E RaMIAs quantify privacy risks more comprehensively than MIAs
@@
As we have explained before, data points that are close enough to the training data are out of the scope of membership inference attacks. We observe from Figure 4 and 10 that range membership inference attacks are better at identifying those nearby points, and thus providing more comprehensive privacy auditing on all the four datasets we tested. We want to highlight that this is the hardest setting for RaMIA, where each IN range only contains one training point that is also not the range center. We also want to emphasize that the gain is remarkable if we consider how little samples were taken compared to the range sizes. On Purchase-100, there are a total of 1024 candidates, and we take less than 20% of them. On AG News, there are millions of sentences within a distance of 8. 50 sentences are too little to meaningfully cover anything in the space. Yet, limited samples can lead to noticeable gains, which further shows the current privacy quantification approach is suboptimal and needs a better framework. Due to randomness in sampling, we report the average gain of RaMIA over MIA with standard deviation in Table III.TPRs at small FPRs are in Table IV. The improvement in AUC is summarized in Table III.
@@
TABLE III: Improvement in AUCs after switching from MIA to RaMIA across multiple iterations of random sampling. We do not randomly sample but use all available attack images in CelebA.
@@
Note that in the CelebA experiment, we use identify information as the semantic range function, making it similar to user-level inference. This further shows that RaMIA is a better and more comprehensive privacy auditing framework. In terms of algorithms, our attack strategy should dominate the simple averaging approach used in prior work [27, 28], since the trimming ratio is optimized (compared to no trimming in simple averaging). The comparison can be found in Fig 11(b). There are other user-level inference algorithms [29] that train shadow models, which incurs additional computational costs. But these methods usually consider scenarios where user information is the label, e.g. facial and speech recognition systems that predict user ID as their outputs [29, 30]. This encourages their target models to explicitly cluster data based on user information, making inference easier. On the other hand, our target model, a facial attribute classifier, does not use user information anywhere in the training, making the inference harder.
TABLE IV: True Positive Rate under different attacks on different datasets at small false positive rates of $1\%$ and $0.1\%$. MIAs cannot be conducted on incomplete data, so we fill the missing columns with the modes and run the attack on them. Standard deviation over random sampling iterations is reported, except for CelebA, where we use all available candidates. The TPR and FPR are calculated based on the range membership information, as described in Sec V-D. For a fair comparison, we should compare RaMIA and MIA based on the same membership testing backbone, e.g. MIA with RMIA versus RaMIA with RMIA. As we have argued, the test queries are unlikely to be exact matches to training points, hence the MIAs are evalauted on range centers, which should be considered non-members. Hence, the TPR values are not indicative of the attack power. This table is only for interested readers who wants to know the attack performance at small FPRs.
@@
LOSS $0\pm 0$ $0\pm 0$ $0.88\pm 0.06$ $0.09\pm 0.03$ 1.40 0.28 $1.10\pm 0.11$ $0\pm 0$
RMIA $2.57\pm 1.58$ $0.57\pm 0.47$ $1.41\pm 0.00$ $0.24\pm 0.00$ 1.44 0.22 $0.54\pm 0.12$ $0\pm 0$
LiRA $2.47\pm 1.47$ $0.50\pm 0.41$ $1.53\pm 0.00$ $0.58\pm 0.00$ 1.10 0.02 $0.63\pm 0.13$ $0\pm 0$
V-F Factors affecting RaMIA performance
@@
Due to the nature of the sampling-based approach, the chance of our attack set including a true training point scales linearly with the density of training points in the range. For a fixed sample size, increasing the range without introducing more training points hurts the attack performance because the chance of the attack set including any training point gets diluted, and vice versa. Figure 5(b) shows that the performance of RaMIA increases when the range becomes larger in the CIFAR-10 experiment. Since the range function in CIFAR-10 is based on image augmentation methods, increasing the range means the attacker applies more distinct augmentation methods, which effectively increases the chance of the attacker obtaining one of the transformed versions of training images seen by the model during training, thus leading to better attack performance. In Figure 4(b), we conduct the attack assuming the attacker cannot sample any true training images. As a sanity check, we relax this assumption, and Figure 5(a) shows that RaMIA performs monotonically better when the density of training images increases from 0% to 50%, when the number of samples is fixed.
(a) RaMIA on CelebA gets better when the training points available for sampling increases. The percentages are the density of training points in the sampling space.
(b) RaMIA performs better on CIFAR-10 when the range size increases. The size is equal to the number of distinct transformations ($n$) applied to images.
@@
(a) Purchase-100 with $\ell_{1}$ range functions
(b) CIFAR-10 with $\ell_{2}$ range functions
Figure 6: Attack performance decreases when increasing range sizes $d$ without introducing more training members to ranges, which effectively reduces the training data density in the range.
@@
We also observe a monotonic decrease in RaMIA performance as the range size increases without introducing true members (Fig 6). We use L1 distance as the range function for Purchase-100, and L2 distance for CIFAR-10, and construct ranges around each data point in MIA’s evaluation set. We ensure that each IN range only contains one training point, which is the range center, and that each OUT range does not contain any training point.
@@
Ranges containing training points that are susceptible to MIAs are also more susceptible to RaMIAs. Researchers have previously discovered that machine learning models memorize duplicate data more [43, 44]. In our CelebA dataset, each celebrity has a different number of photos in the training set, which can be thought of as each identity having different levels of duplication in the training set. Similar to the insights from MIAs, we also observe that identities that have more training images, i.e. higher duplication rate, are more susceptible to RaMIA. Figure 8 shows the relationship between the percentile of each range’s RaMIA score within non-members’ RaMIA scores and the duplication rate. Generally speaking, identities that have more training photos are more prone to RaMIAs. Similarly, correlation can be observed on the other three datasets in our experiments, where the training points’ RaMIA score percentiles among non-members are positively correlated with their MIA score percentiles (Fig 7).
@@
V-G Mismatched training and attack data hurts attack performance
Figure 3(c) shows that MIA underestimates the privacy risk when the augmentation used in training and attacking differs. This should be alarming as many people audit the privacy risk of image classifiers with original images, when the classifiers are often trained with a composition of augmentations. Many transformations, such as color jittering and affine transformations, always produce different final images. Other commonly used augmentation methods, such as random cropping, introduce more randomness to the pipeline. Hence, it is almost certain that the original images are never seen by the model. Therefore, we should use RaMIA for a better auditing result (Figure 4(c)).
V-H RaMIA on redacted data
Many large language models (LLMs) are trained with sensitive textual data. Some of the data with sensitive information redacted might be publicly available. Similar to our experiment with data with missing values, we can apply RaMIA to redacted data to identify which of them are used to train a target LLM. Accurately identifying the redacted sentences paves the way for potentially better data extraction and reconstruction attacks. Figure 9 shows the results. In this experiment, we use spaCy [45] to mask peoples’ names to simulate the masking of personally identifiable information (PII). We then generate 10 possible sentences for each masked sentence using BERT and conduct RaMIA. The MIA performance is the average attack performance over all 10 possible sentences. The performance gap is smaller compared to that in Figure 4(d). The reason might be that BERT fails to produce diverse PII completions, making all candidate sentences similar to each other, and reducing the power of RaMIA.
@@ V-H RaMIA on redacted data
VI Conclusion
We have shown that traditional membership inference attacks (MIAs), which only consider information leakage at exact training points, fail to capture the broader privacy risks in similar but not identical data. We show that when the query shifts from exact training points to nearby points, MIA performance degrades drastically under the correct privacy notion. To address this shortcoming, we introduced range membership inference attacks that evaluates whether a given range contains any training data. Our formulation casts the problem as a composite hypothesis test and proposes a robust trimmed averaging approach to aggregate membership scores over a set of sampled points.
Our proof-of-concept experiments on tabular, image and text datasets demonstrate that RaMIA outperforms MIAs in realistic scenarios with non-exact queries. We acknowledge that RaMIA presents a new technical challenge and that our current formulation and solution are not optimal; rather, they are intended as a starting point to stimulate further research in this promising and meaningful direction. While there is ample room for improving the sampling process and refining the range function design, our work highlights the potential of range-based privacy auditing and motivates future efforts to develop more powerful and robust RaMIA strategies.
We hope that our work can encourage privacy researchers and practitioners to re-examine the conventional MIA paradigm and consider range-based approaches as a more realistic and comprehensive tool for privacy auditing.
Acknowledgement
The authors would like to thank Milad Nasr for the discussion and exploration at the early stage of this project. This research is supported by the Ministry of Education, Singapore, Academic Research Fund (AcRF) Tier 1 (A-8001610-00-00) and Tier 2 (MOE-T2EP20223-0015).